### 27 Jan 2024 v5

# A Novel Skip Orthogonal List for Dynamic Optimal Transport Problem

Abstract
[…]
1 Introduction
The discrete optimal transport (OT) problem involves finding the optimal transport plan “$X$” that minimizes the total cost of transporting one weighted dataset $A$ to another $B$, given a cost matrix “$C$” [23]. The datasets $A$ and $B$ respectively represent the supply and demand node sets, and the problem can be represented as a minimum cost flow problem by adding the edges between $A$ and $B$ to create a complete bipartite graph. The discrete optimal transport problem finds numerous applications in the areas such as image registration [15], seismic tomography [19], and machine learning [31]. However, most of these applications only consider static scenario where the weights of the datasets and the cost matrix remain constant. Yet, many real-world applications need to consider the dynamic scenarios:
  • Dataset Similarity. In data analysis, measuring the similarity between datasets is a crucial task, and optimal transport has emerged as a powerful tool for this purpose [2]. Real-world datasets are often dynamic, with data points being replaced, weights adjusted, or new data points added over time. Therefore, it is necessary to take these dynamically changes into account.
  • Time Series Analysis. Optimal transport can serve as a metric in time series analysis [7]. The main intuition lies in the smooth transition of states between time points in a time series. The smoothness implies the potential to iteratively refine a new solution based on the previous one, circumventing the need for a complete recomputation.
  • Neuroimage analysis [14, 17]. In the medical imaging applications, we may want to compute the change trend of a patient’s organ (e.g., the MRI images of human brain over several months), and the differences are measured by the optimal transport cost. Since the changes are often local and small, we may hope to apply some efficient method to quickly update the cost over the period.
Denote by $V$ and $E$ the sets of vertices and edges in the bipartite network, respectively. Existing methods, such as the Sinkhorn algorithm [9] and the Network Simplex algorithm [21], are not adequate to handle the dynamic scenarios. Upon any modification to the cost matrix, the Sinkhorn algorithm requires at least one Sinkhorn-Knopp iteration to regularize the entire the solution matrix, while the Network Simplex algorithm needs to traverse all edges at least once. Consequently, the time complexities of these algorithms for the dynamic model are $\Omega(\absolutevalue{V}^{2})$ in general cases.
Our algorithm takes a novel data structure that yields an $O(s|V|)$ time solution for handling evolving datasets, where $s$ is determined by the magnitude of the modification. In practice, $s$ is usually much less than the data size $|V|$, and therefore our algorithm can save a large amount of runtime for the dynamic scenarios.  [^1] Demo library is available at https://github.com/xyxu2033/DynamicOptimalTransport
1.1 Related Works
[…]
Approximate Algorithms. For approximate optimal transport, Sherman [25] proposed a $(1+\epsilon)$ approximation algorithm in $\epsilon^{-2}|E|^{1+o(1)}$ time. Pele and Werman [22] introduced the FastEMD algorithm that applies classic algorithms on a heuristic sketch of the input graph. Later, Cuturi [9] used Sinkhorn-Knopp iterations to approximate the optimal transport problem by adding the smoothed entropic entry as the regularization term. Recently several optimizations on the Sinkhorn algorithm have been proposed, such as the Overrelaxation of Sinkhorn  [30] and the Screening Sinkhorn algorithm  [1].
Search Trees and Skip Lists. Our data structure also utilizes high-dimensional extensions of skip lists to maintain a 2-dimensional Euler Tour sequence. Existing high-dimensional data structures based on self-balanced binary search trees, such as $k$-d tree [3], are not suitable as they do not support cyclic ordered set maintenance. Skip lists [24] as depicted in Figure 0(a), which are linked lists with additional layers of pointers for element skipping, is adapted in our context to form skip orthogonal lists. This skipping technique is later generalized for sparse data in higher dimension [20, 12], but range querying generally requires $O(n^{2})$ time where $n^{2}$ is the number of points in the high dimensional space. On the other hand, our data structure requires expected $O(n)$ time when applied to simplex iterations.
1.2 Overview of Our Algorithm
Our algorithm for the dynamic optimal transport problem employs two key strategies:
First, the dynamic optimal transport operations are reduced to simplex iterations. Our technique, grounded on the Simplex method, operates by eliminating the smallest cycle in the graph. We assume that the modifications influence only a small portion of the result, requiring only a few simplex iterations. It is worth noting that existing algorithms like the Network Simplex Algorithm perform poorly under dynamic modifications as they require scanning all the edges at least once to ensure the correctness of the solution.
Second, an efficient data structure is proposed for performing each simplex iteration within expected linear time complexity. Our data structure, as shown in Figure 0(b), employs the Euler Tour Technique. We adapt skip lists to maintain the cyclic ordered set produced by the Euler Tour Technique and introduce an additional dimension to create a Skip Orthogonal List. This structure aids in maintaining the information about the adjusted cost matrix, which is a matrix that requires specific range modifications and queries.
(a) An example for 1D Euler Tour Tree
(b) Our 2D Euler Tour Tree
Figure 1: Overview of Euler Tour Tree with Skip Lists
The rest of the paper is organized as follows. In Section 2, we introduce several important definitions and notations that are used throughout this paper. In Section 3 we present the data structure Skip Orthogonal List, where subsection 3.1 explains how the data structure is organized and subsection 3.2 uses cut operation as an example to demonstrate the updates on this data structure. In Section 4 we elaborate on how to use our data structure to solve the dynamic optimal transport problem. Subsection 4.1 shows that the dynamic optimal transport model can be reduced to simplex iterations, and Subsection 4.2 shows how our data structure could be used to improve the performance of each simplex iteration.
2 Preliminaries
2.1 Optimal Transport
Let $A$ and $B$ represent the source and target point sets, respectively; the corresponding discrete probability distributions are $\alpha\in\mathbb{R}_{\geq 0}^{A}$ and $\beta\in\mathbb{R}_{\geq 0}^{B}$, such that $\sum_{a\in A}\alpha_{a}=\sum_{b\in B}\beta_{b}=1$. The cost matrix is $C\in\mathbb{R}^{A\times B}$ with each entry $c_{ab}$ denoting the cost of transporting a unit of mass from the point $a\in A$ to the point $b\in B$. The discrete optimal transport problem can be formulated as (1).
$\mathcal{W}(\alpha,\beta,C)\triangleq$ $\min_{X\in\mathbb{R}_{\geq 0}^{A\times B}}\sum_{a\in A}\sum_{b\in B}c_{ab}x_{ab}$ (1)
subject to $\begin{cases}\sum_{b\in B}x_{ab}=\alpha_{a}&\forall a\in A\\ \sum_{a\in A}x_{ab}=\beta_{b}&\forall b\in B\end{cases}$
Since Problem (1) is a standard network flow problem, it can be transformed to the following Problem (2) by adding infinity-cost edges [23]:
$\mathcal{W}(w,C)\triangleq$ $\min_{X\in\mathbb{R}_{\geq 0}^{V\times V}}\sum_{u\in V}\sum_{v\in V}c_{uv}x_{uv}$ (2)
subject to $\sum_{v\in V}x_{uv}-\sum_{v\in V}x_{vu}=w_{u}\quad\forall u\in V$
$c_{a_{1}a_{2}}=c_{b_{1}a_{1}}\!\!=c_{b_{1}b_{2}}=+\infty\quad\!\!\!\forall a_{1},a_{2}\in A,b_{1},b_{2}\in B$ (3)
$w_{u}=\begin{cases}\alpha_{a}&\text{if }u=a\in A\\ -\beta_{b}&\text{if }u=b\in B\end{cases}$ (4)
We add the constraint (3), and also redefine the point weights as (4). Note that in (2), the input weight $w$ must always satisfy $\sum_{v\in V}w_{v}=0$; otherwise, the constraints cannot be satisfied.
We use $X$ to denote a given basic solution in the context of using the simplex method to solve the Optimal Transport problem. We notice that the basic variables always form a spanning tree of the complete directed graph with self loops $G(V,V\times V)$ [8]. Let the dual variables be $\pi\in\mathbb{R}^{V}$, satisfying the following constraint:
$x_{uv}\text{ is a basic variable }\implies\pi_{u}-\pi_{v}=c_{uv}.$ (5)
We then define the adjusted cost matrix $C^{\pi}\in\mathbb{R}^{V\times V}$ as $c^{\pi}_{uv}\triangleq c_{uv}+\pi_{v}-\pi_{u}$, where $c^{\pi}_{uv}$ represents the simplex multipliers for the linear program [21].
2.2 Euler Tour Technique
The Euler Tour Technique is a method for representing a tree $T$ as a cyclic ordered set of linear length [29]. Specifically, given a tree $T(V,E_{T})$, we construct a directed graph $D_{T}(V,E_{D})$ with the same vertex set as follows:
  • For each vertex $v\in V_{T}$, add the self-loop $(v,v)$ to $E_{D}$;
  • For each undirected edge $\quantity(u,v)\in E$, add two directed edges $(u,v)$ and $(v,u)$ to $E_{D}$.
[…]
Definition 1 (Euler Tour Representation)
Given a tree $T$, the Euler Tour representation is an arbitrary sequence of Euler Tour of $D_{T}$ represented by edges. That is, $E_{D}$ with circular order induced by the Euler Tour is an Euler Tour representation.
For the rest of the paper, $E_{D}$ denotes the Euler Tour representation of the tree in the context.
Through Definition 1, we can reduce edge linking, edge cutting, sub-tree weight updating and sub-tree weight querying to constant number of element insertion, element deletion, range weight modification and range weight querying on a circular ordered set [29]. We show in Section 4.1 that the dynamic optimal transport can be reduced to the 2D version of these four operations.
2.3 Orthogonal Lists and Skip Lists
Skip Lists are the probabilistic data structures that extend a singly linked list with forward links at various levels, for improving the search, insertion, and deletion operations. Figure 0(a) illustrates an example for skip list. Each level contains a circular linked list, where the list at a higher level is a subset of the list at a lower level and the bottom level contains all the elements. The nodes at the same level have the same color and are linked horizontally. The corresponding elements in adjacent lists are connected by vertical pointers. We apply this skipping technique to circular singly linked lists in our work. Just as most self-balanced binary search trees, Skip Lists support “lazy propagation” techniques, allowing range modifications within $O(\log n)$ time, where $n$ is the sequence length maintained by the tree [27]. This technique is commonly used in dynamic trees for network problems [29].
A $k$-dimensional Orthogonal List has $k$ orthogonal forward links (it is a standard linked list when $k=1$). Orthogonal lists, which can be singly, doubly, or circularly linked, can maintain the information mapped from the Cartesian product of ordered sets, such as sparse tensors [5]. Figure 3 demonstrates an orthogonal list that maintains a $3\times 4$ matrix. Each node has 2 forward links, denoted by row links (red) and column links (blue). The row links connect the elements in each row into a circular linked list horizontally and the column links connect the elements in each column into a circular linked list vertically.
Figure 2: An Orthogonal List Example.
Figure 2: An Orthogonal List Example.
Figure 3: An illustration for Skip Orthogonal List.
3 Skip Orthogonal List
In this section we introduce our novel data structure Skip Orthogonal List. In Section 4, this data structure is used for dynamically updating optimal transport. Formally, with the help of a Skip Orthogonal List, we can maintain a forest $F(V_{F},E_{F})$ with at most two trees, and a matrix $C^{\pi}\in\mathbb{R}^{V_{F}\times V_{F}}$ that supports the following operations. The first two and the last operations are for the case that $F$ contains only one tree; the other two operations are for the case that $F$ has two trees.
  • Cut. Given an undirected edge $\quantity(u,v)$, remove edge $(u,v)$ from the tree, and split it into two disjoint trees. Let the connected component containing $u$ form the vertex set $V_{1}$, and the connected component containing $v$ form the vertex set $V_{2}$.
  • Insert. Add a new node to $F$ that does not connects with any other node. Let the original nodes form the vertex set $V_{1}$, and the new node itself form the vertex set $V_{2}$.
  • Range Update. Given $x$, for each $(u,v)\in V\times V$, update $c^{\pi}_{uv}$ as equation (6).
    $c^{\pi}_{uv}\leftarrow c^{\pi}_{uv}+\begin{cases}0&(u,v)\in V_{1}\times V_{1}\\ -x&(u,v)\in V_{1}\times V_{2}\\ x&(u,v)\in V_{2}\times V_{1}\\ 0&(u,v)\in V_{2}\times V_{2}\end{cases}$ (6)
  • Link. Given a pair $\{u,v\}$, add the edge $(u,v)$ to the forest; connect two disjoint trees into a single tree, if $u$ and $v$ are disconnected.
  • Global Minimum Query. Return the minimum value of $C^{\pi}$ on the tree.
For the remaining of the section, we construct a data structure with the expected $O(\absolutevalue{V}^{2})$ space complexity, where each operation can be done with the expected $O(\absolutevalue{V})$ time. Section 3.1 shows the overall structure of the data structure and how to query in this data structure. Section 3.2 illustrates the cut operation as an example based on this structure. For other operations (linking, insertion, and range updating), we leave them to appendix H.
3.1 The Overall Structure
As shown in Figure 3, a Skip Orthogonal List is a hierarchical collection of Orthogonal lists, where each layer has fewer elements than the one below it, and the elements are evenly spaced out. The bottom layer has all the elements while the top layer has the least. Formally, it can be defined as Definition 2.
Definition 2
Given a parameter $p$ and a cyclic ordered set $S$, a 2D Skip Orthogonal List $\mathcal{L}$ over the set $S$ is an infinite collection of 2 Dimensional Circular Orthogonal Lists =$\mathcal{L}=\{L_{0},L_{1},\cdots\}$, where
  • $\quantity{h(s)}_{s\in S}$ is a set of |$\absolutevalue{S}$ | independent random variables. The distribution is a geometric distribution with parameter $p$
  • For each $\ell\in\mathbb{N}$, let $L_{\ell}$ be an Orthogonal List whose key contains all the elements in $S_{\ell}\times S_{\ell}$, where $S_{\ell}$ is the cyclic sequence formed by $S_{\ell}\triangleq\quantity{s\in S\mid h_{s}\geq\ell}$
Note that for any pair $u,v\in S$, we use $(u,v)$ to denote the corresponding element in $S\times S$; with a slight abuse of notations, we also use “$(u,v)$ at level $\ell$” to denote the corresponding node at the $\ell$-th level in the Skip Orthogonal List. If level $\ell$ is not specified in the context, $(u,v)$ refers to the node at the bottom level.
[…]
Now we augment this data structure to store some additional information for range updating and global minimum query. Before that, the concept “dominate” needs to be adapted to 2D case defined as Definition 3.
Definition 3
For any positive integer $n$, in a Skip Orthogonal List $\mathcal{L}$ over the cyclic ordered set $S$, suppose $(u_{1},v_{1})$ and $(u_{2},v_{2})$ are 2 elements in $S\times S$. We say the node $(u_{1},v_{1})$ dominates $(u_{2},v_{2})$ at level $\ell$ if and only if the following three conditions are all satisfied:
  • $h(u_{1})\geq l$ and $h(v_{1})\geq\ell$;
  • $u_{1}=u_{2}$ or $\max_{u=u_{1}+1}^{u_{2}}h(u)<\ell$;
  • $v_{1}=v_{2}$ or $\max_{v=v_{1}+1}^{v_{2}}h(v)<\ell$.
Here, for any element $s$ in the cyclic ordered set $S$, we use “$s+1$” to denote the successor of $s$ induced by the cyclic order.
To better understand Definition 3, we illustrate the examples in Figure 3(a) and Figure 3(b). In each figure, each blue node dominates itself and all the yellow nodes, while the red nodes dominate every node in the orthogonal list.
(a) Domination in conventional 1D Skip List
(b) Domination in 2D Skip List
Figure 4: Illustrations of Domination in Skip Lists
We now augment the Skip Orthogonal List of Definition 2. For each node $(u,v)$ at orthogonal list $L_{\ell}$, beside the two forward links and two backward links, we add the following attributes:
  • tag: maintains the tag for lazy propagation for all the nodes dominated by it;
  • min_value: maintains the minimum value among all the nodes dominated by it. Note that when $\ell=0$, it stores the original value of $c^{\pi}_{uv}$ following lazy propagation technique. That is, for any node $x$ in the data structure, after each modification and query, the data structure needs to assure
[…]
    =\ = $\min_{(u,v)\in S\times S,x\textit{ dominates }(u,v)}c^{\pi}_{uv}.$
  • min_index: maintains the index corresponding to min_value attribute.
  • child: points to $(u,v)$ at the orthogonal list $L_{\ell-1}$ if $\ell>0$, and it is invalid if $\ell=0$;
  • parent: points to the node that dominates it if $L_{\ell+1}$ is not empty.
3.2 The Update Operation: Cut
In this subsection, we focus on the update operation “Cut” for a 2D Skip Orthogonal List as an example. Figure 4(a) and Figure 4(b) illustrate the basic idea of the cutting process.
(a) Vertical View (some notations in the figure are defined in appendices)
(b) 3D View
Figure 5: Illustrations of a “cut” operation
Taking an undirected edge $(i,j)$ that needs to be cut as the input, the algorithm can be crudely described as follows:
  1. 1.
    Find the two rows and two columns representing the directed edges $(i,j)$ and $(j,i)$ in $E_{D}$, i.e. the transparent nodes in Figure 4(b) and the shaded nodes in Figure 4(a).
  2. 2.
    Push down the tag attribute of all the nodes alongside the rows and columns, i.e. the red nodes and transparent nodes in Figure 4(b). A node $x$ needs to be pushed down, if some changes happen to the nodes dominated by $x$.
  3. 3.
    Cut the rows and columns, warping up the forward links and backward links of points alongside, as illustrated in Figure 4(a). This operation cuts the original Skip Orthogonal List into four smaller lists.
  4. 4.
    Update the min attribute of the remaining nodes whose tag attribute was pushed down in step 2, i.e. the red nodes in Figure 4(b).
  5. 5.
    Return the four smaller lists that were cut out in step 3.
The generalized lazy propagation to our 2D data structure ensures that only the nodes that are “close” to the two rows and columns are modified, and consequently the updating time is guaranteed to be low. Specifically, the expected time complexity is $O(\absolutevalue{V})$ upon each copy of procedure Cut.
4 Our Dynamic Network Simplex Method
The simplex method performs simplex iterations on some initial feasible basis until the optimal solution is obtained. The simplex iterations are used for refining the current solution under the dynamic changes. In each simplex iteration, some variable with negative simplex multiplier is selected for a copy of procedure Pivot, where one common strategy is to pivot in the variable with the smallest simplex multiplier. In Section 4.1 we focus on defining the dynamic optimal transport operations and using simplex iterations to solve this problem, while in Section 4.2 we analyze the details in each simplex iteration. Our method is presented in the context of the conventional Network Simplex algorithm [8, 21].
4.1 Dynamic Optimal Transport Operations
In an Optimal Transport problem, suppose the nodes in node set $V$ are located in some metric space $\mathcal{X}$, e.g., the Euclidean Space $\mathbb{R}^{k}$. The edge cost $c_{uv}$ is usually defined as the (squared) distance between $u$ and $v$ in the space. Let $w\in\mathbb{R}^{V}$ denote the weight vector as defined in the equation (4). A Dynamic Optimal Transport algorithm should support the following four types of update as well as online query:
  • Spatial Position Modification. Select some supply or demand point $v\in V$ and move $v$ to another point $v^{\prime}\in\mathcal{X}$. This update usually results in the modification on an entire row or column in the cost matrix $C\in\mathbb{R}^{A\times B}$.
  • Weight Modification. Select a pair of supply or demand points $u,v\in V$ with some positive number $\delta\in\mathbb{R}_{+}$. Then update $w_{u}\leftarrow w_{u}-\delta$ and $w_{v}\leftarrow w_{v}+\delta$.
  • Point Deletion. Delete a point $v\in V$ with $w_{v}=0$ (before performing deletion, its weight should be already modified to be 00 via the above “weight modification”, due to the requirement of weight balance for OT).
  • Point Insertion. Select a point $v\notin V$; let $w_{v}=0$ and insert $v$ into set $V$ (after the insertion, we can modify its weight from 00 to a specified value via the “weight modification”).
  • Query. Answer the current Optimal Transport plan and the cost.
These updates do not change the overall weights in the supply and demand sets, and thus $\sum_{v\in V}w_{v}\equiv 0$ and a feasible transport plan always exists. Therefore we can reduce these updates to the operations on simplex basis, and we explain the ideas below:
  • Spatial Position Modification. The original optimal solution is primal feasible but not primal optimal, i.e. not dual feasible. We perform the primal simplex method based on the original optimal solution. When moving a point $v$, we first update the cost matrix $C$, the dual variable $\pi$ and the modified cost $C^{\pi}$ to meet the constraint (5). After that, we repeatedly perform the simplex iterations as long as the minimum value of the adjusted cost $C^{\pi}$ is negative.
  • Weight Modification. The original optimal solution is dual feasible but not primal feasible. We perform the dual simplex iterations based on the original optimal solution. Suppose we attempt to decrease $w_{u}$ and increase $w_{v}$ by $\delta$. To implement this, we send $\delta$ amount of flow from $u$ to $v$ in the residual network by the similar manner of the shortest path augmenting method [11]. Specifically, we send the flow through basic variables. If some variable needs to be pivoted out before the required amount of flow is sent, we pivot in the variable with the smallest adjusted cost, and repeat this process.
  • Point Deletion & Point Insertion. As the deleted/inserted point has weight 00 (even if the weight is non-zero, we can first perform the “weight modification” to modify it to be zero), whether inserting or deleting the point does not influence our result. We maintain a node pool keeping all the supply and demand nodes with 0 weight. Each Point Insertion operation takes some point from this pool and move it to the correct spatial location (i.e., insert a new point), while each Point Deletion operation returns a node to the pool.
Our solution updates the optimal transport plan as soon as an update happens, so we can answer the query for the optimal transport plan and value online. If the number of modified nodes is not large, intuitively the optimal transport plan should not change much, and thus we only need to run a small number of simplex iterations to obtain the OT solution. Assume we need to run $s$ simplex iterations, where we assume $s\ll\absolutevalue{V}$ |. Then the time complexity of our algorithm is $O(s\cdot{\rm Time}_{\rm iter})$ with ${\rm Time_{iter}}$ being the time of each simplex iteration.
4.2 The Details for Simplex Iteration
As discussed in Section 4.1, the dynamic operations on OT can be effectively reduced to simplex iterations. In this section, we review the operations used in the conventional network simplex algorithm, and show how to use the data structure designed in Section 3 for maintaining $C^{\pi}$. The conventional network simplex method relies on the simplex method simplified by some graph properties. A (network) simplex iteration contains the following steps:
  1. 1.
    Select Variable to Pivot in. Select the variables with the smallest adjusted cost $C^{\pi}$ to pivot in. Denote by $x_{i_{\rm in}j_{\rm in}}$ the selected one to be pivoted in.
  2. 2.
    Update Primal Solution. Adding the new variable to the current basis forms a cycle. We send the circular flow in the cycle, until some basic variable in the reverse direction runs out of flow, , through Graph Search (e.g. Depth First Search) or Link/Cut Tree [26]. Denote that node as $x_{i_{\rm out}j_{\rm out}}$, which is to be pivoted out.
  3. 3.
    Update Dual Solution. Update the dual variables $\pi$ and modified cost $C^{\pi}$ to meet the constraint (5), as the new basis, because $C^{\pi}$ will soon be queried in the next simplex iteration.
The selecting step performs a query on the data structure on $C^{\pi}$ for the minimum element, and the dual updating performs an update on the data structure. Though the primal updating step can be done within time $O(\log\absolutevalue{V})$ [29], the conventional network simplex maintains $C^{\pi}$ through brute force. That is, the conventional network simplex brutally traverses through all the adjusted costs and selects the minimum, and updates the adjusted cost one by one after the dual solution is updated. This indicates that the time complexity of each simplex iteration is $O(\absolutevalue{V}^{2})$. Our goal is to reduce this complexity; in particular, we aim to maintain $C^{\pi}$ so that it can answer the global minimum query and perform update when the primal basis changes. In the simplex method, when we decide to pivot in the variable $x_{i_{\rm in}j_{\rm in}}$, we update the dual variables as the following equation (7),
$\pi_{u}^{\prime}\leftarrow\pi_{u}+\begin{cases}c_{i_{\rm in}j_{\rm in}}^{\pi}&u\in V_{1}\\ 0&u\in V_{2}\end{cases}$ (7)
where $V_{1}$ is the set of nodes connected to $i_{\rm in}$ and $V_{2}$ is the set of nodes connected to $j_{\rm in}$ after the edge $x_{i_{\rm out}j_{\rm out}}$ is cut; $\pi$ and $\pi^{\prime}$ are respectively the dual solution before and after pivoting, where $\pi_{u}$ and $\pi^{\prime}_{u}$ are the entries corresponding to the node $u$, for each $u\in V$.
Based on the definition of the adjusted cost matrix $C^{\pi}$, the update objective can be formulated as below:
$c^{\pi^{\prime}}_{uv}=c_{uv}^{\pi}+\pi^{\prime}_{u}-\pi^{\prime}_{v}=c^{\pi}_{uv}+\begin{cases}0&(u,v)\in V_{1}\times V_{1}\\ -c^{\pi}_{i_{\rm in}j_{\rm in}}&(u,v)\in V_{1}\times V_{2}\\ c^{\pi}_{i_{\rm in}j_{\rm in}}&(u,v)\in V_{2}\times V_{1}\\ 0&(u,v)\in V_{2}\times V_{2}\end{cases}$
where $C^{\pi}$ is the adjusted cost matrix with regard to the old dual variables $\pi$ while $C^{\pi^{\prime}}$ regards the new dual variables $\pi^{\prime}$. We present the details for updating the adjusted cost matrix in Algorithm 1. Our Skip Orthogonal List presented in Section 3 is capable of performing the operations cut, add and link in $O(\absolutevalue{V})$ time. Therefore we have the following Theorem 4.1.
Theorem 4.1
Each simplex iteration in the conventional network simplex can be completed within expected $O(\absolutevalue{V})$ time.
Algorithm 1 Adjusted Cost Matrix Update
1:Adjusted cost matrix $C^{\pi}$ corresponding to old solution $\pi$, entering variable $x_{i_{\rm in}j_{\rm in}}$, leaving variable $x_{i_{\rm out}j_{\rm out}}$
2:Updated adjusted cost matrix $C^{\pi^{\prime}}$ corresponding to new dual solution $\pi^{\prime}$
3:$t\leftarrow c^{\pi}_{i_{\rm in}j_{\rm in}}$.
4:Cut the edge $\quantity(i_{\rm out},j_{\rm out})$ in $C^{\pi}$ and denote the 4 pieces as $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$.
5:Range update on $C^{\pi}_{12}$ with increasing all the entries by $-t$.
6:Range update on $C^{\pi}_{21}$ with increasing all the entries by $t$.
7:Link the pieces $\{C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}\}$ by the edge $x_{i_{\rm in}j_{\rm in}}$, and obtain the adjusted $C^{\pi^{\prime}}$ as the output.
5 Experiments
All the experimental results are obtained on a server equipped with 512GB main memory of frequency 3200 MHz; the data structures are implemented in C++20 and compiled by G++ 13.1.0 on Ubuntu 22.04.3. The data structures are compiled to shared objects by PyBind11 [16] to be called by Python 3.11.5. Our code uses the Network Simplex library from [4] to obtain an initial feasible flow.
In our experiment, we use the Network Simplex algorithm [21] and Sinkhorn algorithm [9] from the Python Optimal Transport (POT) library [13]. We test our algorithm for both the Spatial Position Modification and Point Insertion scenarios as described in Section 4.1. We take running time to measure their performances.
Datasets. We study the performance of our algorithm on both synthetic and real datasets. For synthetic datasets, we construct a mixture of two Gaussian distributions in $\mathbb{R}^{784}$, where the points of the same Gaussian distribution share the same label. We also use the popular real-world dataset MNIST [18]. We partition the labels into two groups and compute the optimal transport between them.
Setup. We set the Sinkhorn regularization parameter $\lambda$ as $0.1$ and scale the median of the cost matrix $C$ to be 1. We vary the the node size |$\absolutevalue{V}$ | up to $4\times 10^{4}$. For each dataset, we test the static running time of POT on our machine and executed each dynamic operations 100 times to calculate the means of our algorithm. For Spatial Position Modification, we randomly choose some point and add a random Gaussian noise with a variance of 0.5 in each dimension to it. For Point Insertion, we randomly select a point in the dataset that is outside the current OT instance to insert. We perform a query after these updates and compare with the static algorithms implemented in POT.
Result and Analysis. We illustrate our experimental results in Figure 6. In the dynamic scenarios, our algorithm is about 1000 times faster than static Network Simplex algorithm and 10 times faster than Sinkhorn Algorithm when the size |$\absolutevalue{V}$ | reaches 40000, and reveals stable performance in practice. As the number of nodes grow larger, the advantage of our algorithm becomes more significant. This indicates that our algorithm is fast in the case when the number of simplex iterations is not large, as discussed in Section 4.1. Also, the running time of our algorithm is slightly higher than linear trend. Though each simplex iteration in our method is strictly linear in expectation theoretically, this could be influenced by several practical factors, such as the increment in number of simplex iterations, or the decrement in cache hit rate as the node size grows larger.
Figure 6: The ratio of the execution time of our dynamic algorithm to that of the static algorithms.
6 Conclusion and Future Work
In this paper, we propose a dynamic data structure for the traditional network simplex method. With the help of our data structure, the time complexity of the whole pivoting process is $O(\absolutevalue{V})$ in expectation. However, our algorithm lead to several performance issues in practice. First, as our algorithm stores the entire 2D Skip Orthogonal List data structure, it may take relatively high space complexity. Second, as our algorithm is based on linked data structures, the cache hit rate is not high. An interesting future work for improving our implementation is to develop new algorithms and data structures with similar complexity but being more memory friendly.
G Space Complexity of Skip Orthogonal List
[…]
Theorem G.1
The space complexity of a $n\times n$ 2D Skip Orthogonal List with parameter $p$ is $O(n^{2})$ with high probability.
Proof
Denote the number of nodes in the entire Skip Orthogonal List as $N$.
Assume there are $\ell$ levels in the Skip Orthogonal List. Let $m_{i}$ be the number of nodes on the $i$-th level where $1\leq i\leq\ell$, and 00 when $i>\ell$. It is easy to see that
  • $m_{1}=n$;
  • $m_{i}\geq m_{i+1}$ for all $i\geq 1$, i.e., $\quantity{m_{i}}_{i=1}^{\infty}$ is monotonically decreasing;
  • There are $N=\sum_{i=1}^{\infty}m_{i}^{2}$ nodes in the entire Skip Orthogonal List.
If $m_{i}\geq\frac{4}{(1-p)^{2}}\ln n$, we call the $i$-th level a big level; otherwise we call the $i$-th level a small level. By the monotonicity of $\quantity{m_{i}}_{i=1}^{\infty}$, we can see that, there exists a threshold $\ell_{t}$, such that $i\leq\ell_{t}$ if and only if $i$-th level is a big level, i.e., $i\leq\ell_{t}$ if and only if $m_{i}\geq 4{(1-p)^{2}}\ln n$.
Denote $N_{b}\triangleq\sum_{i=1}^{\ell_{t}}m_{i}^{2}$, i.e., the number of nodes in big levels; denote $N_{s}\triangleq\sum_{i=\ell_{t}+1}^{\infty}m_{i}^{2}$, i.e., the number of nodes in small levels. Therefore $N=N_{b}+N_{s}$. We now give their bound respectively with Lemma 1 and Lemma 2.
Lemma 1
$\Pr[N_{b}>\frac{4}{3-2p-p^{2}}n^{2}]\leq\frac{\log_{2/(1+p)}n}{n^{2}}=O\quantity(\frac{1}{n})$
Proof
First we prove that
$\Pr[m_{i+1}>\frac{1+p}{2}m_{i}\ \bigg{|}\ m_{i}\geq\frac{4}{(1-p)^{2}}\ln n]\leq\frac{1}{n^{2}}$
Denote $\quantity{X_{j}}_{j=1}^{m_{i}}$ be the indicator variable of $m_{i}$ events, such that $X_{j}=1$ if and only if the $j$-th row and column of the $i$-th level exists in $i+1$-th level. By the construction of Skip Orthogonal List, we can see that $\quantity{X_{j}}_{j=1}^{m_{i}}$ are independent Bernoulli variables with parameter $p$. Also, by definition, $m_{i+1}=\sum_{j=1}^{m_{i}}X_{j}$ Therefore
  • By linearity of expectation
    $\mathbb{E}\quantity[m_{i+1}]=\sum_{j=1}^{m_{i}}\mathbb{E}\quantity[X_{j}]=pm_{i}$
  • By Hoeffding’s inequality
    $\Pr[m_{i+1}>\frac{1+p}{2}m_{i}]$ $\leq\exp(-\frac{2\quantity(\frac{1+p}{2}m_{i}-\mathbb{E}\quantity[m_{i+1}])^{2}}{m_{i}})=\exp(-\frac{(1-p)^{2}}{2}m_{i})$
Therefore,
$\Pr[m_{i+1}>\frac{1+p}{2}m_{i}\ \bigg{|}\ m_{i}\geq\frac{4}{(1-p)^{2}}\ln n]\leq\exp(-\frac{(1-p)^{2}}{2}\cdot\frac{4}{(1-p)^{2}}\ln n)=\frac{1}{n^{2}}$
When $\frac{1+p}{2}m_{i}\geq m_{i+1}$ for all $11\leq i\leq\ell_{t}-$ 1,
$N_{b}\sum_{i=1}^{\ell_{t}}m_{i}^{2}\leq\sum_{i=0}^{\infty}m_{1}^{2}\quantity(\frac{1+p}{2})^{2i}=\frac{4}{3-2p-p^{2}}n^{2}$
Therefore, when $N_{b}>\frac{4}{3-2p-p^{2}}n^{2}$, there must exists some $i$ in range $1\leq i\leq\min(\log_{2/(1+p)}n,\ell_{t}-1)$ such that $m_{i+1}>\frac{1+p}{2}m_{i}$, since $m_{1},m_{2},\cdots,m_{\ell_{t}}$ are positive integers. By union bound, this happens with probability less than or equal to $\frac{\log_{2/(1+p)}n}{n^{2}}$.
Lemma 2
$\Pr[N_{s}>\frac{8}{\ln\frac{1}{p}(1-p)^{2}}\ln^{3}n]\leq\frac{4\ln n}{(1-p)^{2}n^{2}}$
Proof
In order to bound $N_{s}$, we separate $\quantity{m_{i}}_{i=\ell_{t}+1}^{\infty}$ into $\left\lfloor\frac{4}{(1-p)^{2}}\ln n\right\rfloor$ parts: the $j$-th part contains the number of levels of which sizes are $j\times j$. Denote the number of levels in the $j$-th part as $k_{j}$. Thus $N_{s}=\sum_{i=\ell_{t}+1}^{\infty}m_{i}^{2}=\sum_{j=1}^{\left\lfloor\frac{4}{(1-p)^{2}}\ln n\right\rfloor}j^{2}k_{j}$
If $k_{j}>0$, which indicates that there are some consecutive levels in the Skip Orthogonal List that indeed have size $j$. Since for all $i\geq 1$, $\Pr[m_{i}=m_{i+1}]=p^{m_{i}}$ (all Bernoulli variables indicating whether the row/column remains in the upper level turns true), and the event $\quantity{[m_{i}=m_{i+1}]}_{i=1}^{\infty}$ are independent, we can see that $\Pr[k_{j}>\frac{2}{j\ln\frac{1}{p}}\ln n]\leq\frac{1}{n^{2}}$.
Therefore, by union bound, with probability at least $1-\frac{4}{(1-p)^{2}}\ln n$ we have $k_{j}\leq\frac{2}{j\ln\frac{1}{p}}\ln n$ for all $1\leq j\leq\frac{4}{(1-p)^{2}}\ln n$ and thus
$N_{s}=\sum_{j=1}^{\left\lfloor\frac{4}{(1-p)^{2}}\ln n\right\rfloor}j^{2}k_{j}\leq\frac{2}{\ln\frac{1}{p}}\ln n\sum_{j=1}^{\left\lfloor\frac{4}{(1-p)^{2}}\ln n\right\rfloor}j\leq\frac{8}{\ln\frac{1}{p}(1-p)^{2}}\ln^{3}n$
which is equivalent to the description of Lemma 2.
By Lemma 1, 2 and union bound, as $p\in(0,1)$ is a constant, we can see that there exists some constant $c_{1},c_{2}$ such that $N<c_{1}n^{2}$ with probability at least $1-\frac{c_{2}}{n}$, which finishes the proof of Theorem G.1.
H Operations on Skip Orthogonal List
This section provides detailed algorithm for operations on our Skip Orthogonal List. We use the notation in Figure 4(a) and denote $a^{(i)}$ as the $i^{\rm th}$ ancestor of node $a$.
H.1 Cut
Algorithm 2 Update (Cut)
1:Skip 2D Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to cut
2:Shattered $C^{\pi}$ after cutting the edge $\quantity{u,v}$
3:procedure Cut($C^{\pi},u,v$)
4:     Find the 4 intersections of row $(u,v)$, row $(v,u)$ and column $(u,v)$, column $(v,u)$. Denote them as $p_{11},p_{12},p_{21},p_{22}$ as Figure 4(a) demonstrates.
5:     Run 4 copies of procedure PushDownLines to clear the tag attribute of relevant nodes.
6:     Run 4 copies of procedure CutLine to break the entire 2D Orthogonal List $C^{\pi}$ into 4 pieces.
7:     Run 4 copies of Procedure PullUpLines to update the min attributes of affected nodes.
8:end procedure
9:procedure CutLine($C^{\pi},p,q$)
10:     for each $x$ in the Euler Tour of the current sub-tree do
11:         Move $p,q$ in the current direction.
12:         Let $p_{b},q_{b}$ be the backward node in the orthogonal direction and $p_{f},q_{f}$ be the forward node.
13:         Iterative calculate $p_{b}^{(i)},p_{f}^{(i)},q_{b}^{(i)},q_{f}^{(i)}$, i.e., the $i$-th ancestor.
14:         for $i=h_{x},h_{x}-1,\cdots,0$ do
15:              Link $p_{b}^{(i)}$ and $q_{f}^{(i)}$. Update parent attributes.
16:              Link $q_{b}^{(i)}$ and $p_{f}^{(i)}$. Update parent attributes.
17:              Call procedure PushDown to update the min and tag attribute of their children.
18:         end for
19:     end for
20:end procedure
21:procedure PullUpLines($C^{\pi},p$)
22:     for each Direction dir do
23:         for each edge $x$ in the Euler Tour of the current direction do
24:              for $i=1,2,\cdots,h_{x}$ do
25:                  Let $p^{(i)}$ be the $i^{\rm th}$ ancestor of $p$.
26:                  Call procedure PullUp to update the min attribute of $p^{(i)}$ by brutally traversing through its children set.
27:              end for
28:              Let $p$ be the next node in the Euler Tour.
29:         end for
30:     end for
31:end procedure
Algorithm 2 describes the cut update to update the structure while updating some aggregate information (i.e. tag and min attributes). Here are some clarifications:
[…]
    Note that when linking adjacent nodes, we need to update the parent attributes of nodes $q_{f}^{(i)}$ and $p_{f}^{(i)}$ accordingly, and nodes below it accordingly. However, every node is affected at most once during the entire CutLine procedure, and the number of nodes affected are sure to be affected by PullUpLines process.
  • Procedure PushDownLines is capable of pushing down the tag attribute of all nodes alongside the rows and columns of the four nodes, i.e. the red nodes in 4(b).
  • Similar to PushDown, procedure PullUpLines updates the min attribute of the red nodes in 4(b).
  • In procedure PushDown, $q$ is in the children set of $p$ if and only if $p$ is the parent of $q$. Later we will use this concept to analyze the time complexity.
  • Procedure PushDownLines is similar to procedure PullUpLines. Instead of calling procedure PushDown which updates parents’ min attribute from lower level to higher level, PullUpLines calls procedure PullUp which clears the tag attribute from higher level to lower level. It pushes down the row and column of the input node and the row and column next to the given node.
From analyses above, we can now show Theorem H.1.
Theorem H.1
The expected time complexity of procedure Cut is $O(\absolutevalue{V})$.
Proof
From Algorithm 2, each PushDownLines, CutLine, procedure PullUpLines only apply constant number of modifications and queries on children of nodes alongside the 2 cut rows and columns , i.e. the red and transparent nodes and their children in Figure 4(b).
We first calculate the expected number of nodes removed in procedure Cut, i.e. the transparent nodes in Figure 4(b). The removed nodes in each level are in the row $\rho_{uv}$, $\rho_{vu}$ and column $\tau_{uv},\tau_{vu}$ where $\rho_{uv}\triangleq\quantity{\bigl{(}(u,v),e\bigr{)}:e\in E(D_{T})}$ and $\tau_{uv}\triangleq\quantity{\bigl{(}e,(u,v)\bigr{)}:e\in E(D_{T})}$. By the definition of node height, node $\bigl{(}(u,v),e\bigr{)}$, $\bigl{(}e,(u,v)\bigr{)}$, $\bigl{(}(v,u),e\bigr{)}$ and $\bigl{(}e,(v,u)\bigr{)}$ are in less than or equal to $h_{e}+1$ orthogonal lists. Therefore, the total number of removed points is less than or equal to $4\sum_{e\in E(D_{T})}(h_{e}+1)$. Thus the expected number of removed nodes is less than or equal to $4\sum_{e\in E(D_{T})}(\mathbb{E}(h_{e})+1)=\frac{4}{p}n=O(n)$.
Next we calculate the children of nodes whose the dominate set is affected after procedure Cut update takes place, i.e. red nodes and their children after the Cut update in Figure 4(b). Similar to counting transparent nodes, the expected number of red nodes is also less than or equal to $\frac{4}{p}n=O(n)$. Now we count the number of their children.
Recall that in Figure 4(a) we denote the 2 diagonal pieces of the big skip list as $C^{\pi}_{11}$ and $C^{\pi}_{22}$ while we denote the 2 non-diagonal pieces of the big skip list as $C^{\pi}_{21}$ and $C^{\pi}_{12}$. We count the number of red nodes and their children in the 4 pieces after procedure Cut respectively. Denote $g(n)$ as the number of nodes in $C^{\pi}_{11}$ whose parent’s children set is affected, i.e. whose parent is a red node as Figure 4(b) demonstrates, when $n=\absolutevalue{E(D_{T_{1}})}$ |. Then $g(\absolutevalue{E(D_{T_{2}})})$ is that of $C^{\pi}_{22}$, and we know that:
  • With probability $(1-p)^{n}$, the height of all $n$ rows and columns are 00, which indicates that every node in the current skip orthogonal list does not have a parent. Further more, as shown in Figure 6(a), where there is only one row and one column beside the row just cut (the red row and column), therefore $g(n)=2n-1$.
  • For $k=1,\cdots,n$, $h_{n}=h_{n-1}=\cdots=h_{n-k+2}=0$ and $h_{n-k+1}>0$ occurs with probability $p^{k-1}(1-p)$, as shown in Figure 6(b), where the yellow nodes are the position of the cut-affected node in the upper layer and the green nodes denote upper layer nodes whose children set is not affected. Therefore, $n^{2}-(n-k)^{2}\leq 2kn$ nodes of the current level is a child of some cut-affected node.
(a) Current Layer is at Top
(b) Current Layer is Not at Top
Figure 7: Illustration of Modified Nodes at Some Layer
For $k=0,1,\cdots,n$, with probability $p^{k}(1-p)^{n-k}\binom{n}{k}$, the upper level contains $k$ nodes, i.e. contain $g(k)$ nodes which is visited in procedure Cut (denote $g(0)=0$). By the linearity of expectation, the following holds:
$\mathbb{E}[g(n)]$ $=(1-p)^{n}(2n-1)+\sum_{k=1}^{n}p^{k-1}(1-p)(n^{2}-(n-k)^{2})+\sum_{k=0}^{n}\binom{n}{k}p^{k}(1-p)^{n-k}\mathbb{E}[g(k)]$
$\leq 2n+\sum_{k=1}^{\infty}p^{k-1}(1-p)2nk+\sum_{k=0}^{n}\binom{n}{k}p^{k}(1-p)^{n-k}\mathbb{E}[g(k)]$
$=\frac{4-2p}{1-p}n+\sum_{k=0}^{n}\binom{n}{k}p^{k}(1-p)^{n-k}\mathbb{E}[g(k)]$
We can prove by induction that $\mathbb{E}[g(n)]\leq\frac{4-2p}{(1-p)^{2}}n$. Therefore, the expected number of visited nodes in procedure Cut (i.e. red nodes and their children in Figure 4(b)) in the 2 diagonal pieces $C^{\pi}_{11}$ and $c^{\pi}_{12}$ does not exceed $\mathbb{E}[g(\absolutevalue{E(D_{T_{1}})})]+\mathbb{E}[g(\absolutevalue{E(D_{T_{2}})})]\leq\frac{4-2p}{(1-p)^{2}}\absolutevalue{E(D_{T})}=O(\absolutevalue{E(D_{T})})$.
Denote $G(n,m)$ as the number of nodes visited in the cut process in each of the 2 non-diagonal pieces $C^{\pi}_{12}$ and $C^{\pi}_{21}$ (their numbers are equal by symmetry) where |$\absolutevalue{E(D_{T_{1}})}=n$ and |$\absolutevalue{E(D_{T_{2}})}=m$. Similarly, we can show that
$\mathbb{E}[G(n,m)]=$ $\ ((1-p)^{n}+(1-p)^{m}-(1-p)^{n+m})(n+m-1)$
$+\sum_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}p^{i+j-2}(1-p)^{2}(nm-(n-i)(m-j))$
$+\sum_{\begin{subarray}{c}0\leq i\leq n\\ 0\leq j\leq m\end{subarray}}\binom{n}{i}\binom{m}{j}p^{i+j}(1-p)^{n+m-i-j}\times\mathbb{E}[G(i,j)]$
[…]
H.2 Link
Procedure Link is very similar to procedure Cut. It can be described as Algorithm 3.
Algorithm 3 Update (Link)
1:Shattered Skip Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to link
2:$C^{\pi}$ after linking the edge $\quantity{u,v}$ is
3:Push down tag attribute of nodes along row $(u,u)$, row $(v,v)$, column $(u,u)$ and column $(v,v)$ through PushDownLines.
4:Let $h_{uv},h_{vu}$ be 2 random variables subject to geometric distribution.
5:Link row $(u,u)$ and row $(v,v)$; column $(u,u)$ and column $(v,v)$ with 2 rows and columns of height $h_{uv}$ and $h_{vu}$ through procedure LinkLines.
6:Use PushUpLines procedure to push up min attributes of ancestors of nodes in row $(u,u)$, row $(u,v)$, row $(v,v)$, row $(v,u)$ and these columns.
Here, procedure LinkLine function is very similar to procedure CutLine. It creates a series of new node according to their heights $h_{uv}$ and $h_{vu}$ to link 2 Skip Orthogonal List pieces together in one direction. As procedure link behaves almost the same as procedure Cut, making constant number of modifications to nodes along the newly linked line, the time complexity for each copy of procedure Link is also $O(\absolutevalue{V})$.
H.3 Insert
For procedure Insert, we only need to add one basic variable to the basis to make it feasible again. Without loss of generality, Assume we need to add node $v$ to the demand node set. Procedure Insert could be described in Algorithm 4.
Algorithm 4 Update (Insert)
1:2D Skip Orthogonal List $C^{\pi}$, node $v$ to insert
2:$C^{\pi}$ after inserting node $v$
3:$u\leftarrow\arg\max_{u\in V}\pi_{u}-C_{uv}$
4:$\pi_{v}\leftarrow\pi_{u}-C_{uv}$
5:Insert $v$ into set $V$ and insert $x_{uv}$ into set of basic variables $X_{B}$.
Since $\pi_{v}=\pi_{u}-C^{uv}$ and $x_{uv}$ is the newly added basic variable, dual variable $\pi$ satisfies Constraint 5. Further more, for all $\tilde{u}\in V$, we have $\pi_{v}=\pi_{u}-C_{uv}=\max_{u\in V}\pi_{u}-C_{uv}\geq\pi_{\tilde{u}}-C_{\tilde{u}v}$. Therefore $C^{\pi}_{\tilde{u}v}=C_{uv}+\pi_{v}-\pi_{u}\geq 0$, which indicates that simplex multipliers related to the newly added variable is non-negative. If the old basis is primal optimal, the new basis will sure be primal optimal too.
Line 3 in Algorithm 4 requires adding an entire row and column to the Skip Orthogonal List based on the existing basic variables $X_{B}$. together with a copy of procedure Link to update $X_{B}$. The expected time complexity for these operations are all $O(\absolutevalue{V})$. Therefore the expected time complexity of each copy of Algorithm 4 is $O(\absolutevalue{V})$.
H.4 Range Update (Add) and Query
The tag attribute of each node stores the value that should be added to each node dominated by it but haven’t been propagated. The min attribute stores the minimum value and the index of the minimum value of all nodes dominated by it. Similar to the lazy propagation technique:
  • For procedure RangeUpdate to do the range update that adds all elements by val amount in the current Skip Orthogonal List piece, we add the tag attribute and min attribute of all nodes by val on the top layer.
  • For query min to do range query on the minimum value and indices of all elements in the Skip Orthogonal List, we visit every top layer node and find the node with the minimum min attribute.
Therefore, to bound the running time of procedure RangeUpadte and query GlobalMinimum, we only need to give bound to the nodes on the top level. To achieve this, we show Lemma 4 and 5. But in order to prove them, we need to prove Lemma 3 first.
Lemma 3
Let $n$ be a positive integer and $p$ be a real number in range $(0,1)$. For any $n$ independent and identically distributed random variables $\quantity{X_{i}}_{i=1}^{n}$ following geometric distribution with parameter $p$, the expected number of maximum elements is less than $\frac{1}{p}$, and the expected square of the number of the maximum elements is less than $\frac{2}{p^{2}}-\frac{1}{p}$, i.e.,
$\mathbb{E}\quantity[\absolutevalue{\quantity{i:X_{i}=\max_{j=1}^{n}X_{j}}}]<\frac{1}{p}$
$\mathbb{E}\quantity[\absolutevalue{\quantity{i:X_{i}=\max_{J=1}^{n}X_{j}}}^{2}]<\frac{2}{p^{2}}-\frac{1}{p}$
Proof
Let $Y_{i}$ be the number of maximum elements in subsequence of the first $i$ elements, i.e.,
$Y_{i}\triangleq\absolutevalue{\quantity{k:X_{k}=\max_{j=1}^{i}X_{j}}}$ |
Therefore, Lemma 3 is equivalent to $\mathbb{E}\quantity[Y_{n}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{n}]<\frac{2}{p^{2}}-\frac{1}{p}$. We now prove that $\mathbb{E}\quantity[Y_{i}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{i}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$ for all integer $i=1,2,\cdots,n$ by induction.
First, $Y_{1}=1$, because $\quantity{X_{1}}$ contains only 1 element. Therefore, $\mathbb{E}\quantity[Y_{1}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{1}]<\frac{2}{p}\mathbb{E}\quantity[Y_{1}]-\frac{1}{p}$. Next, we prove that $\mathbb{E}\quantity[Y_{i+1}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{i+1}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i+1}]-\frac{1}{p}$ are implied by $\mathbb{E}\quantity[Y_{i}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{i}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$.
When calculating $Y_{i+1}$ from $Y_{i}$, there are 3 cases:
  1. 1.
    $X_{i+1}>\max_{j=1}^{i}X_{j}$. Suppose this happens with probability $P_{1}$. In this case, $Y_{i+1}=1$.
  2. 2.
    $X_{i+1}=\max_{j=1}^{i}X_{j}$. Suppose this happens with probability $P_{2}$. In this case, $Y_{i+1}=Y_{i}+1$.
  3. 3.
    $X_{i+1}<\max_{j=1}^{i}X_{j}$. Suppose this happens with probability $P_{3}$. In this case, $Y_{i+1}=Y_{i}$.
Notice that, by law of total probability,
$P_{1}$ $=\Pr[X_{i+1}>\max_{j=1}^{i}X_{j}]$
$=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}>\ell]$
$=p\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
$P_{2}$ $=\Pr[X_{i+1}=\max_{j=1}^{i}X_{j}]$
$=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}=\ell]$
$=(1-p)\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
Let $t\triangleq\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$. Therefore, $t$ is bounded by the interval $[0,1]$, and we can express $P_{1}=pt$, $P_{2}=(1-p)t$, $P_{1}+P_{2}+P_{3}=1$.
First we prove $\mathbb{E}\quantity[Y_{i+1}]<\frac{1}{p}$. By law of total expectation, we have
$\mathbb{E}\quantity[Y_{i+1}]$ $=1\cdot P_{1}+(\mathbb{E}\quantity[Y_{i}]+1)\cdot P_{2}+\mathbb{E}\quantity[Y_{i}]\cdot P_{3}=\quantity(1-\mathbb{E}\quantity[Y_{i}]p)t+\mathbb{E}\quantity[Y_{i}]$
Because $\mathbb{E}\quantity[Y_{i}]<\frac{1}{p}$ (inductive hypothesis), $01-\mathbb{E}\quantity[Y_{i}]p>$ 0. Since $t\leq 1,0<p<1$, we get the following
$\mathbb{E}\quantity[Y_{i+1}]=\quantity(1-\mathbb{E}\quantity[Y_{i}]p)t+\mathbb{E}\quantity[Y_{i}]\leq 1-\mathbb{E}\quantity[Y_{i}]p+\mathbb{E}\quantity[Y_{i}]=1+(1-p)\mathbb{E}\quantity[Y_{i}]<\frac{1}{p}$
Hence finishes the proof of $\mathbb{E}\quantity[Y_{i+1}]<\frac{1}{p}$.
Next we prove that $\mathbb{E}\quantity[Y_{i+1}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i+1}]-\frac{1}{p}$. Since $Y_{i+1}=(1-\mathbb{E}\quantity[Y_{i}]p)t+\mathbb{E}\quantity[Y_{i}]\geq\mathbb{E}\quantity[Y_{i}]$, by law of total expectation, we have
$\mathbb{E}\quantity[Y_{i+1}^{2}]=$ $\ 1\cdot P_{1}+(\mathbb{E}\quantity[Y_{i}]+1)^{2}\cdot P_{2}+\mathbb{E}^{2}[Y_{i}]\cdot P_{3}$
$=$ $\ \quantity(-\mathbb{E}\quantity[Y_{i}^{2}]p-1+2\mathbb{E}\quantity[Y_{i}])t+2\mathbb{E}\quantity[Y_{i+1}]+\mathbb{E}\quantity[Y_{i}^{2}]-2\mathbb{E}\quantity[Y_{i}]$
Because $\mathbb{E}\quantity[Y_{i}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$, we have $-\mathbb{E}\quantity[Y_{i}^{2}]-1+2\mathbb{E}\quantity[Y_{i}]>0$. Since $t\leq 1,0<p<1$, we get the following
$\mathbb{E}\quantity[Y_{i+1}^{2}]=$ $\ \quantity(-\mathbb{E}\quantity[Y_{i}^{2}]p-1+2\mathbb{E}\quantity[Y_{i}])t+2\mathbb{E}\quantity[Y_{i+1}]+\mathbb{E}\quantity[Y_{i}^{2}]-2\mathbb{E}\quantity[Y_{i}]$
$\leq$ $\ {-\mathbb{E}}\quantity[Y_{i}^{2}]p-1+2\mathbb{E}\quantity[Y_{i}]+2\mathbb{E}\quantity[Y_{i+1}]+\mathbb{E}\quantity[Y_{i}^{2}]-2\mathbb{E}\quantity[Y_{i}]$
$=$ $\ (1-p)\mathbb{E}\quantity[Y_{i}^{2}]-1+2\mathbb{E}\quantity[Y_{i+1}]$
$<$ $\ (1-p)\quantity(\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p})-1+2\mathbb{E}\quantity[Y_{i+1}]$
$\leq$ $\ (1-p)\quantity(\frac{2}{p}\mathbb{E}\quantity[Y_{i+1}]-\frac{1}{p})-1+2\mathbb{E}\quantity[Y_{i+1}]$
$=$ $\ \frac{2}{p}\mathbb{E}\quantity[Y_{i+1}]-\frac{1}{p}$
Therefore, we have finished the proof that $\mathbb{E}\quantity[Y_{n}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$, and therefore $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p^{2}}-\frac{1}{p}$. Thus Lemma 3 is proved.
With Lemma 3, we can now prove Lemma 4, which immediate leads to the expected time complexity upper bound of procedure RangeUpdate.
Lemma 4
If Skip Orthogonal List $C^{\pi}$ of size $n\times n$ is break into 4 pieces $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$ through procedure Cut, then the expected size of the top level of the 2 non-diagonal pieces $C^{\pi}_{12}$ and $C^{\pi}_{21}$ is $O(n)$.
Proof
Suppose constant $p\in(0,1)$ is the parameter of the Skip Orthogonal List. Suppose the 2 non-diagonal pieces have size $m\times k$ and $k\times m$ respectively. By the properties of procedure Cut, $m+k=n-2$.
For the $m\times k$ piece, denote the height variable corresponding to the rows as $\quantity{a_{i}}_{i=1}^{m}$ and the height variable corresponding to the columns as $\quantity{b_{j}}_{j=1}^{k}$. Therefore, $\quantity{a_{i}}_{i=1}^{m}$ and $\quantity{b_{j}}_{j=1}^{k}$ are i.i.d. following geometric distribution with parameter $p$. The height of the entire Skip Orthogonal List is $\min\quantity(\max_{i=1}^{m}a_{i},\max_{j=1}^{k}b_{j})$.
If node $(i^{\prime},j^{\prime})$, i.e., node on the $i^{\prime}$-th row and $j^{\prime}$-th column, is on the top level, then $\min(a_{i^{\prime}},b_{j^{\prime}})=\min\quantity(\max_{i=1}^{m}a_{i},\max_{j=1}^{k}b_{j})$, which is implied by $a_{i^{\prime}}=\max_{i=1}^{m}a_{i}\vee b_{j^{\prime}}=\max_{j=1}^{k}b_{j}$, i.e., $a_{i^{\prime}}$ is the maximum element of $\quantity{a_{i}}_{i=1}^{m}$ or $b_{j^{\prime}}$ is the maximum element of $\quantity{b_{j}}_{j=1}^{k}$.
Let $I_{ij}$ be the indicator variable that node $(i,j)$ is on the top level. Let $I^{(r)}_{i}$ be the indicator variable that $a_{i}$ is the maximum element of $\quantity{a_{i}}_{i=1}^{m}$ and $I^{(c)}_{j}$ be the indicator that $b_{j}$ is the maximum element of $\quantity{b_{j}}_{j=1}^{k}$. By linearity expectation, the number of elements on the top level is $\sum_{i=1}^{m}\sum_{j=1}^{k}\mathbb{E}\quantity[I_{ij}]$, the number of maximum elements of array $\quantity{a_{i}}_{i=1}^{m}$ is $\sum_{i=1}^{m}\mathbb{E}\quantity[I^{(r)}_{i}]$ and the maximum elements of array $\quantity{b_{j}}_{j=1}^{k}$ is $\sum_{j=1}^{k}\mathbb{E}\quantity[I^{(c)}_{j}]$. From the discussion above, by union bound, we have
$\sum_{i=1}^{m}\sum_{j=1}^{k}\mathbb{E}\quantity[I_{ij}]=$ $\ \sum_{i=1}^{m}\sum_{j=1}^{k}\Pr[\min(a_{i},b_{j})=\min\quantity(\max_{i^{\prime}=1}^{m}a_{i^{\prime}},\max_{j^{\prime}=1}^{k}b_{j^{\prime}})]$
$\leq$ $\ \sum_{i=1}^{m}\sum_{j=1}^{k}\Pr[a_{i}=\max_{i^{\prime}=1}^{m}a_{i^{\prime}}\vee b_{j}=\max_{j^{\prime}=1}^{k}b_{j^{\prime}}]$
$\leq$ $\ \sum_{i=1}^{m}\sum_{j=1}^{k}\Pr[a_{i}=\max_{i^{\prime}=1}^{m}a_{i^{\prime}}]+\Pr[b_{j}=\max_{j^{\prime}=1}^{k}b_{j^{\prime}}]$
$\leq$ $\ k\sum_{i=1}^{m}\mathbb{E}\quantity[I^{(r)}_{i}]+m\sum_{j=1}^{k}\mathbb{E}\quantity[I^{(c)}_{j}]$
As $\sum_{i=1}^{m}I^{(r)}_{i}$ equals to the total number of maximum values in $\quantity{a_{i}}_{i=1}^{m}$, by Lemma 3, we have $\sum_{i=1}^{m}I^{(r)}_{i}<\frac{1}{p}$. Similarly, $\sum_{j=1}^{k}\mathbb{E}\quantity[I^{(c)}_{j}]<\frac{1}{p}$. As a result, the total number of variables on the top level is smaller than $k\cdot\frac{1}{p}+m\cdot\frac{1}{p}<\frac{n}{p}=O(n)$, which finishes the proof of Lemma 4.
Further more, with Lemma 3, we can directly derive Lemma 5.
Lemma 5
The expected size of the top level of a Skip Orthogonal List $C^{\pi}$ is $O(1)$.
Now, with Lemma 4 and 5, the time complexity for procedure RangeUpdate and Query are directly obtained.
I Supplementary Experiments
Similar to the previous experiments, the experimental results in this section are obtained from the same server with the same dataset pools, i.e. Synthetic dataset about Gaussian Distributions on $\mathbb{R}^{784}$ and MNIST dataset [18]. Our algorithm is compared with Network Simplex Algorithm [21] and Sinkhorn Algorithm [9] by Python Optimal Transport Library [13]. For test Weight Modification, we randomly select a pair of nodes in the system, and send flow of which amount is in the range of ±$\pm 1\%$ of their node weight difference, repeat for 100 times; for test Point Deletion & Insertion, we first delete 100 nodes with 0 weight in the current data set to return it to the pool, and insert 100 new nodes in the system.
Figure 8: Supplementary Experiments
Figure 8 demonstrates our experiment result, where the time ratio of our algorithm and their algorithm are plotted. From the figure, the asymptotic advantage of our algorithm is obvious. where $T_{\rm our}$ is the running time of each dynamic operation of our algorithm, while $T_{\rm net}$ and $T_{\rm sinkhorn}$ are the static running time of Network Simplex algorithm and Sinkhorn algorithm respectively. Our algorithm for handling point insertion only takes less than $1\%$ of time compared to static algorithms when |$\absolutevalue{V}$ | reaches $40000$. Also the Insert and Delete procedures perform much more stable than space position modification and weight modification procedures, regardless of which dataset we use. This matches the theory that procedure Insert and Delete is strictly linear, while the performance of normal space position modification and weight modification procedures depends on the number of simplex iterations it experiences. Though procedure Insert and Delete must be combined with weight modification procedures to obtain the result, Figure 8 shows that procedure Insert and Delete are not a great overhead.